# Quantum teleportation implies symmetry-protected topological order

### v2

@@ 2.2 Quantum state transfer
Both tasks transfer $k\geq 1$ “logical states” within a given many-body state. A logical state is simply a particular state of a quantum bit—i.e., $\lvert\psi_{n}\rangle=\alpha_{n}\,\lvert 0\rangle_{n}+\beta_{n}\lvert 1\rangle_{n}$ for the $n$th “logical qubit” with logical basis states $\lvert 0\rangle_{n}$ and $\lvert 1\rangle_{n}$. The simplest protocols $\mathcal{W}$ transfer logical states between individual physical qubits; however, logical states may also be “embedded” in many physical qubits.
Logical qubits can also be characterized in terms of their associated logical operators. If the many-body state $\lvert\Psi\rangle$ encodes $k\geq 1$ logical states $\lvert\psi_{n}\rangle$ with coefficients $\alpha_{n},\beta_{n}$ (for $1\leq n\leq k$), then the corresponding logical basis operators $X^{(n)}_{{\rm L}},Y^{(n)}_{{\rm L}},Z^{(n)}_{{\rm L}}$ are defined such that,
*$\left\langle*\middle|\Psi\middle|X^{(n)}_{{\rm L}}\right\rangle{\Psi}$ =*$=\left\langle*\middle|\psi_{n}\middle|X\right\rangle{\psi_{n}}=2\,\mathrm{Re}(\alpha^{*}_{n}\beta_{n})$ (4a)
*$\left\langle*\middle|\Psi\middle|Y^{(n)}_{{\rm L}}\right\rangle{\Psi}$ =*$=\left\langle*\middle|\psi_{n}\middle|Y\right\rangle{\psi_{n}}=2\,\mathrm{Im}(\beta^{*}_{n}\alpha_{n})$ (4b)
*$\left\langle*\middle|\Psi\middle|Z^{(n)}_{{\rm L}}\right\rangle{\Psi}$ =*$=\left\langle*\middle|\psi_{n}\middle|Z\right\rangle{\psi_{n}}=\lvert\alpha_{n}\rvert^{2}-\lvert\beta_{n}\rvert^{2}\,,~{}~{}$ (4c)
@@
In general, the logical operators must reproduce the Pauli algebra and satisfy Eq. 2.2. However, in the context of state transfer and teleportation, we assume that the initial and final many-body states are separable with respect to the $k$ logical qubits and all other physical qubits, and realized on $k$ physical sites. Often, the $n$th logical qubit is realized on the physical qubit $j$, so that $X^{(n)}_{{\rm L}},Y^{(n)}_{{\rm L}},Z^{(n)}_{{\rm L}}$ reduce to the standard Paulis $X_{j}$, $Y_{j}$, and $Z_{j}$.
@@
A unitary protocol $\mathcal{W}$ acting on $\mathcal{H}_{\rm ph}$ (1) achieves quantum state transfer of $k$ logical qubits from the initial logical vertex set $I=\{i_{1},\dots,i_{k}\}\subset V$ to the final logical vertex set $F=\{f_{1},\dots,f_{k}\}\subset V$ if
$\lvert\Psi_{T}\left(\boldsymbol{\alpha},\boldsymbol{\beta}\right)\rangle=\mathcal{W}\,\lvert\Psi_{0}\left(\boldsymbol{\alpha},\boldsymbol{\beta}\right)\rangle\,,~{}~{}$ (5)
@@
$\lvert\Psi_{0}\left(\boldsymbol{\alpha},\boldsymbol{\beta}\right)\rangle$ $=\left[\bigotimes\limits_{n=1}^{k}\,\lvert\psi_{n}(\alpha_{n},\beta_{n})\rangle_{i_{n}}\right]\otimes\lvert\Phi_{0}\rangle_{I^{c}}$ (6)
$\lvert\Psi_{T}\left(\boldsymbol{\alpha},\boldsymbol{\beta}\right)\rangle$ $=\left[\bigotimes\limits_{n=1}^{k}\,\lvert\psi_{n}(\alpha_{n},\beta_{n})\rangle_{f_{n}}\right]\otimes\lvert\Phi_{T}\rangle_{F^{c}}\,,~{}$ (7)
where the logical state $\lvert\psi_{n}\rangle=\alpha_{n}\lvert 0\rangle+\beta_{n}\lvert 1\rangle$ is transferred from site $i_{n}\in I$ (6) to site $f_{n}\in F$ (7), $\boldsymbol{\alpha}=(\alpha_{1},\dots,\alpha_{k})$ and $\boldsymbol{\beta}=(\beta_{1},\dots,\beta_{k})$ satisfy $\lvert\alpha_{n}\rvert^{2}+\lvert\beta_{n}\rvert^{2}=1$ for each $n$, and $\lvert\Phi_{0}\rangle$ and $\lvert\Phi_{T}\rangle$ are arbitrary many-body states on the $N-k$ qubits in $I^{c}$ and $F^{c}$, respectively.
We now present Prop. 1, which establishes necessary and sufficient conditions for quantum state transfer in terms of logical and stabilizer operators, an equivalence between the transfer of logical states and logical operators, and that the logical operators in the initial and final many-body states are spanned by the Pauli operators $X_{i_{n}},Z_{i_{n}}$ and $X_{f_{n}},Z_{f_{n}}$, for $i_{n}\in I$ and $f_{n}\in F$, respectively.
@@
Given an initial state of the form (6), the following conditions are necessary and sufficient for a unitary protocol $\mathcal{W}$ to achieve quantum state transfer of $k$ logical states $\lvert\psi_{n}\rangle$,
$\mathcal{W}^{\dagger}X_{f_{n}}\mathcal{W}$ $=X_{i_{n}}\,\mathsf{S}_{n,x}$ (8a)
$\mathcal{W}^{\dagger}Z_{f_{n}}\mathcal{W}$ $=Z_{i_{n}}\,\mathsf{S}_{n,z}\,,$ (8b)
for $1\leq n\leq k$, where $\mathsf{S}_{n,\nu}\in\operatorname*{stab}\left(\lvert\Psi_{0}\rangle\right)$ (3) are unitary and stabilize $\lvert\Psi_{0}\rangle$ (6); see also Def. 1.
The proof of Prop. 1 appears in App. A. Loosely speaking, Prop. 1 establishes that the task of transferring $k$ logical states from the logical sites $I=\{i_{1},\dots,i_{k}\}\subset V$ to the logical sites $F=\{f_{1},\dots,f_{k}\}\subset V$ in the Schrödinger picture is equivalent to transferring $k$ pairs of $X$- and $Z$-type logical operators from the logical sites in $F$ to those in $I$ in the Heisenberg picture. The result for the $n$th $Y$-type logical operator follows from $Y=\mathrm{i}XZ$:
$\mathcal{W}^{\dagger}Y_{f_{n}}\mathcal{W}$ $=Y_{i_{n}}\,\mathsf{S}_{n,y}\,,$ (9)
where $\mathsf{S}_{n,y}\equiv Z_{i_{n}}\mathsf{S}_{n,x}Z_{i_{n}}\mathsf{S}_{n,z}\in\operatorname*{stab}\left(\lvert\Psi_{0}\rangle\right)$.
Importantly the result (1) of Prop. 1 captures how any logical operation $\mathcal{O}_{n}$ on the $n$th logical qubit in the final state $\lvert\Psi_{T}\rangle$ (7) is related to the same logical operation $\mathcal{O}_{n}$ on the $n$th logical qubit in the initial state $\lvert\Psi_{0}\rangle$ (6). Since the logical Paulis $\sigma^{\nu}_{n,{\rm L}}$ for $\nu=1,2,3$ form a complete basis for $\mathcal{O}_{n}$ in the final state $\lvert\Psi_{T}\rangle$ (7), Prop. 1 implies that
$O_{f_{n}}(T)$ $=\mathcal{W}^{\dagger}\,O_{f_{n}}\mathcal{W}$
$=\frac{1}{2}\sum\limits_{\nu=0}^{3}\,\operatorname*{tr}\limits_{f_{n}}\left[O_{f_{n}}\,\sigma^{\nu}_{f_{n}}\right]\,\mathcal{W}^{\dagger}\,\sigma^{\nu}_{f_{n}}\,\mathcal{W}$
$=\frac{1}{2}\sum\limits_{\nu=1}^{3}\,\operatorname*{tr}\limits_{i_{n}}\left[O_{i_{n}}\,\sigma^{\nu}_{i_{n}}\right]\,\sigma^{\nu}_{i_{n}}\,\mathsf{S}_{n,\nu}$
$\equiv O_{i_{n}}\mathsf{S}_{n,O}\,,$ (10)
is equivalent to $O_{i_{n}}$ acting on $\lvert\Psi_{0}\rangle$ (6).
@@
which is magnitude of the largest eigenvalue of $A$. Using the fact that $\lVert\left\{\sigma^{\nu}_{n,{\rm L}},\,\mathsf{S}\right\}\rVert=2$, we find that Friedman et al. (2022a)
$\lVert\left[\mathcal{W}^{\dagger}X_{f_{n}}\mathcal{W},\,Z_{i_{n}}\right]\rVert=2\,,$ (12)
meaning that successful state transfer implies that the Heisenberg-evolved logical operator $X_{f_{n}}(T)$ for the $n$th logical qubit realizes the operator $X_{i_{n}}$ acting on the initial state, and obeys the Pauli algebra (12).
@@
The task distance (or teleportation distance) for a protocol $\mathcal{W}$ that maps the initial state $\lvert\Psi_{0}\rangle$ (6) with $k$ logical qubits on the sites $I=\{i_{1},i_{2},\dots,i_{k}\}\subset V$ to the state $\lvert\Psi_{T}\rangle$ (7) with logical qubits on the sites $F=\{f_{1},f_{2},\dots,f_{k}\}\subset V$ is given by
@@ 2.3 Stinespring measurement
To facilitate our discussion of quantum teleportation—which involves measurements and outcome-dependent feedback—we now review the Stinespring formalism Friedman et al. (2022a, b); Barberena and Friedman (2023); Stinespring (1955), which derives from the Stinespring Dilation Theorem Stinespring (1955); Choi (1975); Takesaki (2012). The formalism leads to a unitary representation of projective measurements and outcome-dependent operations in a dilated Hilbert space,
$\mathcal{H}_{\rm dil}\equiv\mathcal{H}_{\rm ph}\otimes\mathcal{H}_{\rm ss}\,,~{}~{}$ (14)
where $\mathcal{H}_{\rm ph}$ is the physical Hilbert space (1) and $\mathcal{H}_{\rm ss}$ is the “Stinespring” Hilbert space. For a protocol $\mathcal{W}$ involving $M$ single-qubit measurements, $\mathcal{H}_{\rm ss}=\mathbb{C}^{2M}$ contains $M$ qubits, which store the $M$ outcomes (and reflect the states of the corresponding measurement apparati).
The crucial result of the Stinespring Theorem Stinespring (1955); Takesaki (2012); Choi (1975) in this context is that all valid quantum channels (i.e., positive maps) can be represented via isometries. An isometry is a map $\mathbb{V}\,:\,\mathcal{H}\to\mathcal{H}^{\prime}$, where $\mathcal{D}\leq\mathcal{D}^{\prime}$. When $\mathcal{D}=\mathcal{D}^{\prime}$, $\mathbb{V}$ is a unitary; when $\mathcal{D}<\mathcal{D}^{\prime}$, $\mathbb{V}$ dilates $\mathcal{H}$ to $\mathcal{H}^{\prime}=\mathcal{H}_{\rm dil}$ (14). Crucially, one can always embed the isometry $\mathbb{V}\,:\,\mathcal{H}_{\rm ph}\to\mathcal{H}_{\rm dil}$ in a unitary $U\,:\,\mathcal{H}_{\rm dil}\to\mathcal{H}_{\rm dil}$ by working in $\mathcal{H}_{\rm dil}$ from the outset; this merely requires identifying a default initial state of all measurement apparati, which we take to be $\lvert 0\rangle_{\rm ss}$ without loss of generality Takesaki (2012); Choi (1975); Friedman et al. (2022b, a); Stinespring (1955); Barberena and Friedman (2023).
We first consider the Stinespring representation of projective measurements. An observable $A$ with $\mathcal{N}$ unique eigenvalues $a_{n}$ has a spectral decomposition given by
$A$ $=\sum\limits_{n=0}^{\mathcal{N}-1}\,a_{n}\,P^{(n)}_{A}\,,~{}~{}$ (15)
@@
$\mathcal{M}_{\left[A\right]}$ $=\sum\limits_{m,n=0}^{\mathcal{N}-1}\,P^{(m)}_{A}\otimes\lvert n+m\rangle\langle n\rvert_{\rm ss}\,,~{}~{}$ (16)
which is unique up to the definition of the Stinespring basis and choice of default state Stinespring (1955); Takesaki (2012); Friedman et al. (2022b, a); Barberena and Friedman (2023); Choi (1975). Projecting the post-measurement state onto the $n$th Stinespring state leads to the physical state $\lvert\psi^{\prime}\rangle\propto P^{(n)}_{A}\,\lvert\psi\rangle$ (15), as required. For example, when $A=Z_{j}$, we have
$\mathcal{M}_{\left[Z_{j}\right]}$ $=\frac{1}{2}\sum\limits_{m=0,1}\left(\mathds{1}+(-1)^{m}\,Z_{j}\right)\otimes\widetilde{X}^{m}_{j}~{}~{}$
$=P^{(0)}_{j}\otimes\widetilde{\mathds{1}}_{j}+P^{(1)}_{j}\otimes\widetilde{X}_{j},$ (17)
@@
$\mathcal{R}_{\Omega}$ $=\sum\limits_{\boldsymbol{n}}\,\chi_{\boldsymbol{n}}\otimes\widetilde{P}^{(\boldsymbol{n})}_{\Omega}\,,~{}~{}$ (18)
where $\widetilde{P}^{(\boldsymbol{n})}_{\Omega}$ is an $\widetilde{Z}$-basis projector onto the outcome “trajectory” $\boldsymbol{n}=(n_{1},\dots,n_{s})$ for the subset $\Omega\subset V_{\rm ss}$. The “recovery operators” $\chi_{\boldsymbol{n}}$ must act unitarily on $\mathcal{H}_{\rm ph}$ and/or any unused Stinespring registers. For example, one might apply $X_{i}$ to the physical qubit $i$ if the $j$th outcome was $-1$, and do nothing otherwise, so that $\mathcal{R}_{\Omega}$ (18) becomes
$\mathcal{R}_{j\to i}$ $=\mathds{1}_{i}\otimes\widetilde{P}^{(0)}_{j}+X_{i}\otimes\widetilde{P}^{(1)}_{j}\,,~{}~{}$ (19)
which is simply a CNOT gate from the Stinespring site $j$ to the physical qubit $i$, in contrast to $\mathcal{M}_{[Z]}$(2.3).

### v3

@@ 2.2 Quantum state transfer
Both tasks transfer $k\geq 1$ “logical states” within a given many-body state. A logical state is simply a particular state of a quantum bit—i.e., $\lvert\psi^{\,}_{n}\rangle=\alpha^{\,}_{n}\,\lvert 0\rangle^{\,}_{n}+\beta^{\,}_{n}\lvert 1\rangle^{\,}_{n}$ for the $n$th “logical qubit” with logical basis states $\lvert 0\rangle^{\,}_{n}$ and $\lvert 1\rangle^{\,}_{n}$. The simplest protocols $\mathcal{W}$ transfer logical states between individual physical qubits; however, logical states may also be “embedded” in many physical qubits.
Logical qubits can also be characterized in terms of their associated logical operators. If the many-body state $\lvert\Psi\rangle$ encodes $k\geq 1$ logical states $\lvert\psi^{\,}_{n}\rangle$ with coefficients $\alpha^{\,}_{n},\beta^{\,}_{n}$ (for $1\leq n\leq k$), then the corresponding logical basis operators $X^{(n)}_{{\rm L}},Y^{(n)}_{{\rm L}},Z^{(n)}_{{\rm L}}$ are defined such that,
$\left\langle*\middle|\Psi\middle|X^{(n)}_{{\rm L}}\right\rangle{\Psi}$ $=\left\langle*\middle|\psi^{\vphantom{\prime}}_{n}\middle|X^{\vphantom{*}}\right\rangle{\psi^{\vphantom{\prime}}_{n}}=2\,\mathrm{Re}(\alpha^{*}_{n}\beta^{\vphantom{\prime}}_{n})$ (4a)
$\left\langle*\middle|\Psi\middle|Y^{(n)}_{{\rm L}}\right\rangle{\Psi}$ $=\left\langle*\middle|\psi^{\vphantom{\prime}}_{n}\middle|Y^{\vphantom{*}}\right\rangle{\psi^{\vphantom{\prime}}_{n}}=2\,\mathrm{Im}(\beta^{*}_{n}\alpha^{\vphantom{\prime}}_{n})$ (4b)
$\left\langle*\middle|\Psi\middle|Z^{(n)}_{{\rm L}}\right\rangle{\Psi}$ $=\left\langle*\middle|\psi^{\vphantom{\prime}}_{n}\middle|Z^{\vphantom{*}}\right\rangle{\psi^{\vphantom{\prime}}_{n}}=\lvert\alpha^{\vphantom{\prime}}_{n}\rvert^{2}-\lvert\beta^{\vphantom{\prime}}_{n}\rvert^{2}\,,~{}~{}$ (4c)
@@
In general, the logical operators must reproduce the Pauli algebra and satisfy Eq. 4. However, in the context of state transfer and teleportation, we assume that the initial and final many-body states are separable with respect to the $k$ logical qubits and all other physical qubits, and realized on $k$ physical sites. Often, the $n$th logical qubit is realized on the physical qubit $j$, so that $X^{(n)}_{{\rm L}},Y^{(n)}_{{\rm L}},Z^{(n)}_{{\rm L}}$ reduce to the standard Paulis $X^{\vphantom{*}}_{j}$, $Y^{\vphantom{*}}_{j}$, and $Z^{\vphantom{*}}_{j}$.
@@
A unitary protocol $\mathcal{W}$ acting on $\mathcal{H}^{\,}_{\rm ph}$ (1) achieves quantum state transfer of $k$ logical qubits from the initial logical vertex set $I=\{i_{1},\dots,i_{k}\}\subset V$ to the final logical vertex set $F=\{f_{1},\dots,f_{k}\}\subset V$ if
$\lvert\Psi^{\vphantom{\prime}}_{T}\left(\boldsymbol{\alpha},\boldsymbol{\beta}\right)\rangle=\mathcal{W}\,\lvert\Psi^{\vphantom{\prime}}_{0}\left(\boldsymbol{\alpha},\boldsymbol{\beta}\right)\rangle\,,~{}~{}$ (5)
@@
$\lvert\Psi^{\vphantom{\prime}}_{0}\left(\boldsymbol{\alpha},\boldsymbol{\beta}\right)\rangle$ $=\left[\bigotimes\limits_{n=1}^{k}\,\lvert\psi^{\vphantom{*}}_{n}(\alpha^{\vphantom{*}}_{n},\beta^{\vphantom{*}}_{n})\rangle^{\vphantom{\dagger}}_{i_{n}}\right]\otimes\lvert\Phi^{\vphantom{\prime}}_{0}\rangle^{\vphantom{\dagger}}_{I^{c}}$ (6)
$\lvert\Psi^{\vphantom{\prime}}_{T}\left(\boldsymbol{\alpha},\boldsymbol{\beta}\right)\rangle$ $=\left[\bigotimes\limits_{n=1}^{k}\,\lvert\psi^{\vphantom{*}}_{n}(\alpha^{\vphantom{*}}_{n},\beta^{\vphantom{*}}_{n})\rangle^{\vphantom{\dagger}}_{f_{n}}\right]\otimes\lvert\Phi^{\vphantom{\prime}}_{T}\rangle^{\vphantom{\dagger}}_{F^{c}}\,,~{}$ (7)
where the logical state $\lvert\psi^{\,}_{n}\rangle=\alpha^{\,}_{n}\lvert 0\rangle+\beta^{\,}_{n}\lvert 1\rangle$ is transferred from site $i^{\,}_{n}\in I$ (6) to site $f^{\,}_{n}\in F$ (7), $\boldsymbol{\alpha}=(\alpha^{\,}_{1},\dots,\alpha^{\,}_{k})$ and $\boldsymbol{\beta}=(\beta^{\,}_{1},\dots,\beta^{\,}_{k})$ satisfy $\lvert\alpha^{\,}_{n}\rvert^{2}+\lvert\beta^{\,}_{n}\rvert^{2}=1$ for each $n$, and $\lvert\Phi^{\,}_{0}\rangle$ and $\lvert\Phi^{\,}_{T}\rangle$ are arbitrary many-body states on the $N-k$ qubits in $I^{c}$ and $F^{c}$, respectively.
We now present Prop. 1, which establishes necessary and sufficient conditions for quantum state transfer in terms of logical and stabilizer operators, an equivalence between the transfer of logical states and logical operators, and that the logical operators in the initial and final many-body states are spanned by the Pauli operators $X^{\vphantom{*}}_{i_{n}},Z^{\vphantom{*}}_{i_{n}}$ and $X^{\vphantom{*}}_{f_{n}},Z^{\vphantom{*}}_{f_{n}}$, for $i^{\,}_{n}\in I$ and $f^{\,}_{n}\in F$, respectively.
@@
Given an initial state of the form (6), the following conditions are necessary and sufficient for a unitary protocol $\mathcal{W}$ to achieve quantum state transfer of $k$ logical states $\lvert\psi^{\,}_{n}\rangle$,
$\mathcal{W}^{\dagger}X^{\vphantom{*}}_{f_{n}}\mathcal{W}$ $=X^{\vphantom{*}}_{i_{n}}\,\mathsf{S}_{n,x}$ (8a)
$\mathcal{W}^{\dagger}Z^{\vphantom{*}}_{f_{n}}\mathcal{W}$ $=Z^{\vphantom{*}}_{i_{n}}\,\mathsf{S}_{n,z}\,,$ (8b)
for $1\leq n\leq k$, where $\mathsf{S}_{n,\nu}\in\operatorname*{stab}\left(\lvert\Psi^{\,}_{0}\rangle\right)$ (3) are unitary and stabilize $\lvert\Psi^{\,}_{0}\rangle$ (6); see also Def. 1.
The proof of Prop. 1 appears in App. A. Loosely speaking, Prop. 1 establishes that the task of transferring $k$ logical states from the logical sites $I=\{i_{1},\dots,i_{k}\}\subset V$ to the logical sites $F=\{f_{1},\dots,f_{k}\}\subset V$ in the Schrödinger picture is equivalent to transferring $k$ pairs of $X$- and $Z$-type logical operators from the logical sites in $F$ to those in $I$ in the Heisenberg picture. The result for the $n$th $Y$-type logical operator follows from $Y^{\vphantom{*}}=\mathrm{i}X^{\vphantom{*}}Z^{\vphantom{*}}$:
$\mathcal{W}^{\dagger}Y^{\vphantom{*}}_{f_{n}}\mathcal{W}$ $=Y^{\vphantom{*}}_{i_{n}}\,\mathsf{S}_{n,y}\,,$ (9)
where $\mathsf{S}_{n,y}\equiv Z^{\vphantom{*}}_{i_{n}}\mathsf{S}_{n,x}Z^{\vphantom{*}}_{i_{n}}\mathsf{S}_{n,z}\in\operatorname*{stab}\left(\lvert\Psi^{\vphantom{*}}_{0}\rangle\right)$.
Importantly the result (8) of Prop. 1 captures how any logical operation $\mathcal{O}^{\,}_{n}$ on the $n$th logical qubit in the final state $\lvert\Psi^{\,}_{T}\rangle$ (7) is related to the same logical operation $\mathcal{O}^{\,}_{n}$ on the $n$th logical qubit in the initial state $\lvert\Psi^{\,}_{0}\rangle$ (6). Since the logical Paulis $\sigma^{\nu}_{n,{\rm L}}$ for $\nu=1,2,3$ form a complete basis for $\mathcal{O}^{\,}_{n}$ in the final state $\lvert\Psi^{\,}_{T}\rangle$ (7), Prop. 1 implies that
$O^{\vphantom{\prime}}_{f_{n}}(T)$ $=\mathcal{W}^{\dagger}\,O^{\vphantom{\prime}}_{f_{n}}\mathcal{W}$
$=\frac{1}{2}\sum\limits_{\nu=0}^{3}\,\operatorname*{tr}\limits_{f_{n}}\left[O^{\vphantom{\prime}}_{f_{n}}\,\sigma^{\nu}_{f_{n}}\right]\,\mathcal{W}^{\dagger}\,\sigma^{\nu}_{f_{n}}\,\mathcal{W}$
$=\frac{1}{2}\sum\limits_{\nu=1}^{3}\,\operatorname*{tr}\limits_{i_{n}}\left[O^{\vphantom{\prime}}_{i_{n}}\,\sigma^{\nu}_{i_{n}}\right]\,\sigma^{\nu}_{i_{n}}\,\mathsf{S}_{n,\nu}$
$\equiv O^{\vphantom{\prime}}_{i_{n}}\mathsf{S}_{n,O}\,,$ (10)
is equivalent to $O^{\,}_{i_{n}}$ acting on $\lvert\Psi^{\,}_{0}\rangle$ (6).
@@
which is magnitude of the largest eigenvalue of $A$. Using the fact that $\lVert\left\{\sigma^{\nu}_{n,{\rm L}},\,\mathsf{S}\right\}\rVert=2$, we find that [42]
$\lVert\left[\mathcal{W}^{\dagger}X^{\vphantom{*}}_{f_{n}}\mathcal{W},\,Z^{\vphantom{*}}_{i_{n}}\right]\rVert=2\,,$ (12)
meaning that successful state transfer implies that the Heisenberg-evolved logical operator $X^{\vphantom{*}}_{f_{n}}(T)$ for the $n$th logical qubit realizes the operator $X^{\vphantom{*}}_{i_{n}}$ acting on the initial state, and obeys the Pauli algebra (12).
@@
The task distance (or teleportation distance) for a protocol $\mathcal{W}$ that maps the initial state $\lvert\Psi^{\,}_{0}\rangle$ (6) with $k$ logical qubits on the sites $I=\{i_{1},i_{2},\dots,i_{k}\}\subset V$ to the state $\lvert\Psi^{\,}_{T}\rangle$ (7) with logical qubits on the sites $F=\{f_{1},f_{2},\dots,f_{k}\}\subset V$ is given by
@@ 2.3 Stinespring measurement
To facilitate our discussion of quantum teleportation—which involves measurements and outcome-dependent feedback—we now review the Stinespring formalism [42, 47, 46, 48], which derives from the Stinespring Dilation Theorem [48, 49, 50]. The formalism leads to a unitary representation of projective measurements and outcome-dependent operations in a dilated Hilbert space,
$\mathcal{H}^{\vphantom{\dagger}}_{\rm dil}\equiv\mathcal{H}^{\vphantom{\dagger}}_{\rm ph}\otimes\mathcal{H}^{\vphantom{\dagger}}_{\rm ss}\,,~{}~{}$ (14)
where $\mathcal{H}^{\,}_{\rm ph}$ is the physical Hilbert space (1) and $\mathcal{H}^{\,}_{\rm ss}$ is the “Stinespring” Hilbert space. For a protocol $\mathcal{W}$ involving $M$ single-qubit measurements, $\mathcal{H}^{\,}_{\rm ss}=\mathbb{C}^{2M}$ contains $M$ qubits, which store the $M$ outcomes (and reflect the states of the corresponding measurement apparati).
The crucial result of the Stinespring Theorem [48, 50, 49] in this context is that all valid quantum channels (i.e., positive maps) can be represented via isometries. An isometry is a map $\mathbb{V}\,:\,\mathcal{H}\to\mathcal{H}^{\prime}$, where $\mathcal{D}\leq\mathcal{D}^{\prime}$. When $\mathcal{D}=\mathcal{D}^{\prime}$, $\mathbb{V}$ is a unitary; when $\mathcal{D}<\mathcal{D}^{\prime}$, $\mathbb{V}$ dilates $\mathcal{H}$ to $\mathcal{H}^{\prime}=\mathcal{H}^{\,}_{\rm dil}$ (14). Crucially, one can always embed the isometry $\mathbb{V}\,:\,\mathcal{H}^{\,}_{\rm ph}\to\mathcal{H}^{\,}_{\rm dil}$ in a unitary $U\,:\,\mathcal{H}^{\,}_{\rm dil}\to\mathcal{H}^{\,}_{\rm dil}$ by working in $\mathcal{H}^{\,}_{\rm dil}$ from the outset; this merely requires identifying a default initial state of all measurement apparati, which we take to be $\lvert 0\rangle^{\,}_{\rm ss}$ without loss of generality [50, 49, 47, 42, 48, 46].
We first consider the Stinespring representation of projective measurements. An observable $A$ with $\mathcal{N}$ unique eigenvalues $a^{\,}_{n}$ has a spectral decomposition given by
$A$ $=\sum\limits_{n=0}^{\mathcal{N}-1}\,a^{\vphantom{\prime}}_{n}\,P^{(n)}_{A}\,,~{}~{}$ (15)
@@
$\mathcal{M}^{\vphantom{\dagger}}_{\left[A\right]}$ $=\sum\limits_{m,n=0}^{\mathcal{N}-1}\,P^{(m)}_{A}\otimes\lvert n+m\rangle\langle n\rvert^{\vphantom{\prime}}_{\rm ss}\,,~{}~{}$ (16)
which is unique up to the definition of the Stinespring basis and choice of default state [48, 50, 47, 42, 46, 49]. Projecting the post-measurement state onto the $n$th Stinespring state leads to the physical state $\lvert\psi^{\prime}\rangle\propto P^{(n)}_{A}\,\lvert\psi\rangle$ (15), as required. For example, when $A=Z^{\vphantom{*}}_{j}$, we have
$\mathcal{M}^{\vphantom{\dagger}}_{\left[Z^{\vphantom{*}}_{j}\right]}$ $=\frac{1}{2}\sum\limits_{m=0,1}\left(\mathds{1}+(-1)^{m}\,Z^{\vphantom{*}}_{j}\right)\otimes\widetilde{X}^{m}_{j}~{}~{}$
$=P^{(0)}_{j}\otimes\widetilde{\mathds{1}}^{\vphantom{*}}_{j}+P^{(1)}_{j}\otimes\widetilde{X}^{\,}_{j},$ (17)
@@
$\mathcal{R}^{\vphantom{\dagger}}_{\Omega}$ $=\sum\limits_{\boldsymbol{n}}\,\chi^{\vphantom{\dagger}}_{\boldsymbol{n}}\otimes\widetilde{P}^{(\boldsymbol{n})}_{\Omega}\,,~{}~{}$ (18)
where $\widetilde{P}^{(\boldsymbol{n})}_{\Omega}$ is an $\widetilde{Z}^{\vphantom{*}}$-basis projector onto the outcome “trajectory” $\boldsymbol{n}=(n^{\,}_{1},\dots,n^{\,}_{s})$ for the subset $\Omega\subset V^{\,}_{\rm ss}$. The “recovery operators” $\chi^{\,}_{\boldsymbol{n}}$ must act unitarily on $\mathcal{H}^{\,}_{\rm ph}$ and/or any unused Stinespring registers. For example, one might apply $X^{\vphantom{*}}_{i}$ to the physical qubit $i$ if the $j$th outcome was $-1$, and do nothing otherwise, so that $\mathcal{R}^{\,}_{\Omega}$ (18) becomes
$\mathcal{R}^{\vphantom{\dagger}}_{j\to i}$ $=\mathds{1}^{\vphantom{\dagger}}_{i}\otimes\widetilde{P}^{(0)}_{j}+X^{\vphantom{*}}_{i}\otimes\widetilde{P}^{(1)}_{j}\,,~{}~{}$ (19)
which is simply a CNOT gate from the Stinespring site $j$ to the physical qubit $i$, in contrast to $\mathcal{M}^{\,}_{[Z]}$(2.3).